OpenStudy (anonymous):

What is the value of w? 5(9 − 3) = w • 9 – 5 • 3

imqwerty (imqwerty):

1 Simplify 9−3 to 6 5×6=w×9−5×3 2 Simplify 5×6 to 30 30=w×9−5×3 3 Regroup terms 30=9w−5×3 4 Simplify 5×3 to 15 30=9w−15 5 Add 15 to both sides 30+15=9w 6 Simplify 30+15 to 45 45=9w 7 Divide both sides by 9 45/9=w 8 Simplify 459 to 5 5=w

OpenStudy (anonymous):

i really get it now

imqwerty (imqwerty):

:)